### v1

# Chaos controlled and disorder driven phase transitions by breaking permutation symmetry

Abstract
Introducing disorder in a system typically breaks symmetries and can introduce dramatic changes in its properties such as localization. At the same time, the clean system can have distinct many-body features depending on how chaotic it is. In this work the effect of permutation symmetry breaking by disorder is studied in a system which has a controllable and deterministic regular to chaotic transition. Results indicate a continuous phase transition from an area-law to a volume-law entangled phase irrespective of whether there is chaos or not, as the strength of the disorder is increased. The critical disorder strength obtained by finite size scaling, indicate a strong dependence on whether the clean system is regular or chaotic to begin with. In the process, we also obtain the critical exponents associated with this phase transition. Additionally, we find that a relatively small disorder is seen to be sufficient to delocalize a chaotic system.
Deterministic chaos, originating in the lack of sufficient number of conserved quantities, gives rise to ergodic and mixing dynamics with a positive Lyapunov exponent in the classical case Tabor (1989); Lichtenberg and Lieberman (2013); Ott (2002), and in the quantum case results in states with random matrix properties, thermalization and volume law entanglement Gutzwiller (1991); Stöckmann (1999); Mehta (2004); Wimberger (2014); D’Alessio et al. (2016); Kaufman et al. (2016); Bohigas et al. (1984); Srednicki (1994). Disorder, depending on its relative strength on the other hand, can lead to Griffiths phases Griffiths (1969), Anderson Anderson (1958) or many-body localization Basko et al. (2006); Oganesyan and Huse (2007), and can also result in thermalization and delocalization via breaking symmetries. For example, the Sachdev-Ye-Kitaev (SYK) model is one where randomized all-to-all couplings results in maximal quantum chaos, maximizing a temperature dependent quantum Lyapunov exponent. Sachdev and Ye (1993); Kitaev ; Maldacena and Stanford (2016).
Arrays of superconducting Josephson junction based transmons, which are nonlinear oscillators, are currently the most successful implementations of quantum computers Koch et al. (2007); Barends et al. (2013); Wang et al. (2022). Disorder in the energies of the individual qubit transmons is a necessary evil in the presence of inter-transmon couplings. A detailed study of the consequences for these devices comes with a warning of being “dangerously close to a phase of uncontrollable chaotic fluctuations” Berke et al. (2022).
On the other hand, it has long been appreciated that dynamical chaos has serious effects on practical implementation of quantum many-body systems, as it can increase entanglement to random state values, and small subsystems face effective decoherence, thus questioning the reliability of quantum information processing Zurek (2006); Hauke et al. (2012). Several related aspects have been addressed in Refs. Bandyopadhyay and Lakshminarayan (2002); Gong and Brumer (2003); Flambaum (2000); Song and Shepelyansky (2001); Georgeot and Shepelyansky (2000); Braun (2002); Madhok et al. (2018); Shepelyansky (2001); Piga et al. (2019); Chinni et al. (2021); Berke et al. (2022); Lerose and Pappalardi (2020), for example it has been discussed in Ref. Georgeot and Shepelyansky (2000) that the presence of random individual qubit energies along with residual short range interactions can make the system chaotic under certain limits that compromises the quantum computer. In another work, it is shown that the error in quantum computation may increase exponentially with system size in the chaotic regime in the presence of imperfections Song and Shepelyansky (2001). Thus, it is important to understand the interplay between disorder and dynamical chaos. We therefore, in this work, bring together the clearly connected, yet disparate, fields of chaos and disorder, and study the effect of disorder on regular or chaotic quantum dynamics.
In particular, we investigate the dynamics of $N$ particles originally restricted to the small $O(N)$ permutation symmetric subspace (PSS), but that can have various dynamical regimes, from near-integrable to intermediate and fully chaotic. The disorder breaks the permutation symmetry and results in a phase transition into the full $O(\exp(N))$ Hilbert space. Interestingly, the extent of chaos in the “clean” symmetric space is found to dictate the critical strength of the disorder that ensures the dynamics has penetrated to the full Hilbert space, in the large $N$ limit. The transition has many possible order parameters. The block entanglement changes from area law $\sim\ln N$, characteristic of permutation symmetric states Seshadri et al. (2018) to volume law $\sim N$. Collective observables such as the total angular momentum changes from $\sim N$ to $\sim\sqrt{N}$. This transition shows similarities to a transition from the superradiant to the radiant phases of collective atomic states in a cavity Scully (2015); Gegg et al. (2018); Rubies-Bigorda and Yelin (2022). A recent work studied the dynamics when the system does not start from a permutation symmetric subspace of an otherwise permutation symmetric Hamiltonian Qi et al. (2023). They find non-generic effects of the deep Hilbert space by studying growth of out-of-time order correlator and entanglement. Another work studied the effect of introducing non-uniform initial states on the permutation symmetry of the Hamiltonian Iemini et al. (2024). However, we wish to study the case when the disorder free system starting from PSS has a controlled transition to deterministic chaos, so that its interplay with disorder becomes transparent.
A convenient model for this study is provided by the kicked top which is a textbook example of quantum chaos Haake et al. (1987); D’Ariano et al. (1992); Haake (2013), of one large spin of angular momentum $S=N/2$, that can be considered as the collective dynamics of $N$ spin $1/2$ particles within the permutation symmetric subspace Wang et al. (2004); Madhok et al. (2015). Several experiments have fruitfully implemented this to understand the effects of chaos on entanglement Chaudhury et al. (2009); Neill et al. (2016); Krithika et al. (2019) . A disordered version of the kicked top can be used to investigate the effects of breaking permutation symmetry. The specific Hamiltonian that we study is given by
$H=\frac{k}{\tau 2N}\sum_{\ell<\ell^{\prime}=1}^{N}(1+\epsilon_{\ell\ell^{\prime}})\sigma_{\ell}^{x}\sigma_{\ell^{\prime}}^{x}+\frac{p}{2}\sum_{n=-\infty}^{\infty}\sum_{\ell=1}^{N}\sigma_{\ell}^{y}\delta(t-n\tau),$ (1)
where, $\sigma_{\ell}^{\nu}$ with $\nu=x~{},y,~{}z$ are the Pauli matrices at site $\ell$, and $k$ is the interaction strength that can be tuned such that there is a transition from regular to deterministic chaotic dynamics in the disorder free Hamiltonian, $\tau$ is the time period of delta kicks that is set to unity, and $p$ is fixed to $\pi/2$ to compare with the known results. Here, ϵsubscriptitalic-$\epsilon_{\ell\ell^{\prime}}$ are the random (quenched disorder) part of the spin interactions taken from a normal distribution with zero mean and standard deviation $w$, also referred to as disorder strength. The corresponding unitary Floquet operator $U_{w}$ that evolves states infinitesimally before consecutive kicks separated by the period $\tau$, is given by
$\exp\left({-i\frac{k}{2N}\sum_{\ell<\ell^{\prime}=1}^{N}(1+\epsilon_{\ell\ell^{\prime}})\sigma_{\ell}^{x}\sigma_{\ell^{\prime}}^{x}}\right)\exp\left({-i\frac{\pi}{4}\sum_{\ell=1}^{N}\sigma_{\ell}^{y}}\right),\$ (2)
where we have set $\hbar=1$.
The disorder free version $U_{0}$, with all $\epsilon_{\ell\ell^{\prime}}=0$, is the quantum kicked top Haake (2013); Wimberger (2014); Haake et al. (1987); Zyczkowski (1990); Sanders and Milburn (1989); Bhosale and Santhanam (2018); Fox and Elston (1994); Wang et al. (2010); Zou and Wang (2022); Muñoz-Arias et al. (2021); Fiderer and Braun (2018); Bhosale and Santhanam (2017); Sieberer et al. (2019); Lerose and Pappalardi (2020); Kumari and Ghose (2018); Lombardi and Matzkin (2011); Anand et al. (2024) where the Hamiltonian is purely a function of the collective spin operators $J_{\alpha}=\sum_{\ell}^{N}\sigma_{\ell}^{\alpha}/2$ where $\alpha=x,~{}y,~{}z$. The total angular momentum $J^{2}=J_{x}^{2}+J_{y}^{2}+J_{z}^{2}$ defined in terms of these collective spins is a constant of motion in the quantum kicked top Haake (2013); Wang et al. (2004); Ghose et al. (2008). The large $N$ limit of $U_{0}$ is studied as a symplectic classical map on the unit sphere $(X,Y,Z)\mapsto(X^{\prime},Y^{\prime},Z^{\prime})$ where $X=J_{x}/J$, $Y=J_{y}/J$ and $Z=J_{z}/J$. This map shows regular dynamics for small values of $k$, typically $<2$, and fully chaotic dynamics for large values of $k$ ( $k>6$), where the Lyapunov exponent may be approximated as ($\ln k-1$) Constantoudis and Theodorakopoulos (1997); Wang and Robnik (2021). Various aspects of long range interacting systems such as given above are discussed in Ref. Defenu et al. (2024).
If the initial state is an unentangled spin coherent state given by Glauber and Haake (1976); Lee Loh and Kim (2015); Puri (2001)
|$\ket{\psi_{0}}=\ket{\theta,\phi}=\left(\cos\frac{\theta}{2}\ket{0}+e^{i\phi}\sin\frac{\theta}{2}\ket{1}\right)^{\otimes{N}},$ (3)
then the state after $n$ Floquet periods is |$\ket{\psi_{n}}=U_{0}^{n}\ket{\psi_{0}}$. The states at all times $n$ are restricted to the $N+1$ dimensional PSS and conserve the total angular momentum $J^{2}=(N/2)(N/2+1)$. It has also been shown that the chaotic phase can be described using a random matrix theory (RMT) in $N+1$ dimensions Haake (2013). For small values of $N(\leq 4)$, the kicked top is exactly solvable and does yield interesting insights already into the large $N$ cases Ruebeck et al. (2017); Dogra et al. (2019); Sreeram et al. (2021).
Introducing disorder in the system will break the permutation symmetry of the Hamiltonian, and the state $|\psi_{n}\rangle$ after $n$ kicks given by $U_{w}^{n}\ket{\psi_{0}}$ will now span the full $2^{N}$ dimensional Hilbert space where $J^{2}$ is no longer a constant of motion.
Figure 1: Schematic of the dynamics: Top row corresponds to disorder free system where interactions connecting different spins have same magnitude as depicted by connecting arrows of black color. The second term of the Hamiltonian is the periodic kicks along $y-$direction. Bottom row corresponds to a disordered system as shown by connecting arrows of different colors. For times $n\tau<t<(n+1)\tau-\Delta\tau$, the interactions are dominant whereas for $(n+1)\tau-\Delta\tau<t<(n+1)\tau$, it is the field term which is dominant. Here, we use delta kicks which correspond to $\Delta\tau\to 0$ limit. If the radius of the sphere on the right is proportional to $N$, then $\langle J\rangle$ for the disorder free system is restricted on the surface of this sphere, whereas for the disordered system it is pushed inside the sphere where $\langle J\rangle\sim\sqrt{N}$, as also shown in the figure above.
The state |$\ket{\psi_{n}}$ in presence of disorder tends to a random state on the full Hilbert space, denoted as |$\ket{\psi_{\text{RMT}}}$, in which the expectation value of $J_{z}^{2}$ can be estimated by:
$\left\langle\ J_{z}^{2}\right\rangle_{{\text{RMT}}}$ $=$ $\sum_{j=1}^{2^{N}}\lambda_{j}^{2}\overline{|\left\langle\psi_{\text{RMT}}|\phi_{j}\right\rangle|^{2}}=N/4,$ (4)
where |$\ket{\phi_{j}}=\ket{j_{1}\cdots j_{N}}$, $j_{i}=\pm 1$, are eigenvectors of $J_{z}$ with eigenvalues $\lambda_{j}=(\sum_{\ell}j_{\ell})/2$. The ensemble average over random states $\overline{|\left\langle\psi_{\text{RMT}}|\phi_{j}\right\rangle|^{2}}=1/2^{N}$ is used above along with the fact that $\operatorname{Tr}J_{z}^{2}=N2^{N-1}/2$. Similarly $\left\langle J_{x}^{2}\right\rangle_{\text{RMT}}=\left\langle J_{y}^{2}\right\rangle_{\text{RMT}}=N/4$, and hence $\left\langle J^{2}\right\rangle_{\text{RMT}}=3N/4$. Clearly, ${\langle J^{2}\rangle}$ exhibit a change in the scaling when the dynamics shifts from the permutation symmetric subspace where $\langle J^{2}\rangle\sim N^{2}$ to the full Hilbert space where the state being close to random has a $\sim N$ scaling. It is this transition of the dynamics to the “full” Hilbert space that we shall focus in the current work. A schematic diagram of the Hamiltonian studied highlighting the difference in the dynamics of disorder free as well as the disordered system is shown in Fig. 1.
One can also capture this change in the dynamics by studying von Neumann entropy $S_{Q}(A)$ of the subsystem $A$ consisting of $Q-$ spins or qubits. The von Neumann entropy is given by $S_{Q}(A)=-\mbox{tr}(\rho_{A}\log_{2}\rho_{A})$, where $\rho_{A}$ is the reduced density matrix of subsystem $A$. The expression for $S_{Q}(A)$ using ensemble averaging over random permutation symmetric states is approximated by Seshadri et al. (2018)
$\left\langle S_{Q}(A)\right\rangle_{PSS}\approx\log_{2}(Q+1)-\frac{2}{3}\left(\frac{Q+1}{N-Q+1}\right).$ (5)
We have checked numerically that $S_{Q}(A)$ even in the regular regime of small $k$ within permutation symmetric subspace is proportional to ${}\log_{2}(Q+1)$ dependence. A related work calculating the von Neumann entropy in PSS is given in Ref.Castro-Alvaredo and Doyon (2013). On the other hand, when the dynamics traverse the full Hilbert space, $S_{Q}(A)$ is given by the Page value Page (1993); Seshadri et al. (2018); Lee et al. (2017):
$\left\langle S_{Q}(A)\right\rangle_{2^{Q}}\approx Q-\frac{1}{\ln 2}\left(\frac{2^{Q}}{2^{N-Q+1}}\right).$ (6)
As $w$, the standard deviation of the disorder distribution or the disorder strength is increased, various expectation values will potentially shift from their values within the permutation symmetric subspace to the RMT values relevant to the full Hilbert space. This is because increasing the strength of the disorder $w$ will not only take the system out of the permutation symmetric subspace but will also make it more chaotic since increasing $w$ is equivalent to increasing $k$ as can be seen from Eq. 1. Thus as $w\to 0$ or small when the dynamics is predominantly within the permutation symmetric subspace, the scaling of long time limit of various quantities will follow the PSS scaling, $i.e.,$ ${\langle J^{2}\rangle}\sim N^{2}$ and ${\langle{S_{Q}(A)}\rangle}\sim\log_{2}(Q+1)$. In the other extreme limit, when $w\to\infty$ and the dynamics is described by RMT in the full $2^{N}-$dimensions, ${\langle J^{2}\rangle}=3N/4$ and ${\langle{S_{Q}(A)}\rangle}\sim Q$.
We present the results of numerical calculations which are performed with an initial state as given in Eq. 3 with $\theta=2.25,~{}\phi=1.1$ and is evolved using the Floquet operator $U_{w}$ given in Eq. 2 . We use the fast Walsh Hadamard transform to go to system sizes upto $N=16$. The evolution of $\langle J^{2}\rangle$ and $S_{N/2}$ as a function of number of kicks, or time $n$, is shown in Fig. 2 for $k=1$. We find that for small values of $w$, $\langle J^{2}\rangle$ remains a constant till $n\sim N+1$ corresponding to the Heisenberg time relevant to the PSS beyond which the system is taken out of PSS so that its value starts decreasing. It is around $n\sim 2^{N}$, the Heisenberg time relevant to the full Hilbert space that $\langle J^{2}\rangle$ again starts saturating, which implies that the system has realized the presence of full Hilbert space around this time. For all further calculations, we perform time averaging from $n=10^{5}$ to $n=3\times 10^{5}$ when a thermalization has set in for $N=16$, the maximum system size considered in this work. The change in scaling of $J^{2}$ and $S_{N/2}(A)$ as discussed in the text is presented in Fig. 3(a) and (b), respectively, for different $w$.
Figure 2: (a) $\left\langle J^{2}\right\rangle$ with respect to time or number of kicks $n$ for different disorder strengths $w$ with $k=1$, $N=12$ qubits, averaged over 100 disorder realizations. Inset shows the entanglement $S_{N/2}$ as a function of $n$ for the same parameters as in the main figure.
Figure 3: (a) Change in scaling of $\langle J^{2}\rangle$ from $N^{2}/4+N/2$ to $43N/$ 4 as $w$ is increased where the chaos parameter $k$ is set to 1.(b) Same for $S_{N/2}(A)$ as a function of $N$ at $k=1$ as discussed in the text. The red dashed line in Fig. (a) corresponds to $N^{2}/4+N/2$ whereas the black dotted line corresponds to the RMT result $43N/$ 4. Similarly, red dashed line in Fig. (b) corresponds to fitted line $0.41+0.54\log_{2}(N/2+1)$ and the black dotted line is the RMT expression given by $N/2-1/2\ln 2$.
It is tempting to associate this change in the scaling of $\langle J^{2}\rangle$ and $S_{N/2}$ with a second order phase transition. Therefore, we now investigate the possible existence of a critical $w_{c}$ below which the dynamics is predominantly within the permutation symmetric subspace and above which the dynamics captures the properties of the full Hilbert space described by RMT, and obtain relevant critical exponents of the associated phase transition, if any. We therefore propose a finite size scaling of the form Botet and Jullien (1983); Skinner et al. (2019); Lee et al. (2017); Li et al. (2018); Cardy (1988)
${\langle J^{2}\rangle}=N^{\zeta/\nu}F((w-w_{c})N^{1/\nu}),$ (7)
where $\nu$ is the correlation length exponent. Fig. 4 (main) shows the crossing of the data around $w_{c}=2.4$ when $k=1$ whereas the collapse of the data with $\nu=0.59$ and $\zeta=0.59$ is shown in the inset of Fig. 4. In order to confirm this phase transition further, we repeat the calculations for other values of $k$, the results of which are shown in Figure 5. It is to be noted that $k=4$ is a mixed phase space with considerable chaos already present in the system.
Figure 4: $\langle J^{2}\rangle/N^{\zeta/\nu}$ as a function of disorder strength $w$ for $k=1$ and different system sizes. These curves cross each other at $w=w_{c}\approx 2.4$. The inset shows the collapse of the data with $\zeta=0.59$ and $\nu=0.59$ consistent with the scaling proposed in Eq. 7.
Figure 5: (a) The main figure shows the data corresponding to different system sizes for $k=2$ crossing each other at $w=w_{c}\sim 0.41$ whereas the inset shows the collapse of the data with $\nu=0.40$ and $\zeta=0.49$.(b) Same as (a) but for $k=4$ with $w=w_{c}\sim 0.17$ and $\nu=0.3$.
Table 1 lists the critical disorder strength $w_{c}$ along with the critical exponent $\nu$ for different values of the interaction strength or chaos parameter $k$. Intuitively, the chaotic dynamics at larger $k$ values will require smaller disorder to take the dynamics to the full Hilbert space, which is also what we observe. Interestingly, we also find a weak parameter dependence on the critical exponent $\nu$ for this phase transition. The exponent $\nu$ lies within a range of 0.6 to 0.3 for the parameters studied. Such a variation seems to be a characteristic of a disordered system such as the well studied many-body-localization one Wang et al. (2021). We also find that in most cases, $\nu\sim\zeta$. The quality of the data collapse for different values of $k$ confirm the presence of a second order phase transition. At the same time, we do observe that identifying correct $w_{c}$ values for larger $k$ is difficult since $w_{c}$ then is very small and a single crossing point for all system sizes is difficult to obtain. On the other hand when $k\rightarrow 0$, the dynamics can get highly non-ergodic and these parameters need more careful examinations. We have further confirmed this phase transition using the entanglement entropy $S_{N/2}(A)$ with a similar finite size scaling as in Eq. 7 giving $\nu=0.6$, $w_{c}=1.8$ when $k=1$, as shown in Fig. 6.
It is also observed that the collapse is relatively good when the clean system is either predominantly chaotic or predominantly regular. For a mixed phase space, with significant fraction of both regular and chaotic regions, say at $k=3$, the quality of the scaling seems to depend on the initial state. The value in the table 1 for $k=3$ and $k=4$ corresponds to chaotic regions. For regular regions in these cases, we could not observe scaling behaviour. This may be attributed to the small values of angular quantum number $j$, that we are forced to restrict ourselves to, as well as the complexity of mixed phase space (see for example Madhok (2015) for cases when the initial state in a kicked top is at the border between chaos and regular regions) and the influence of phenomena such as chaos assisted tunneling Tomsovic and Ullmo (1994); S.Keshavamurthy and P.Schlagheck ; Krithika et al. (2023), that warrants a detailed separate study.
$k$ $w_{c}$ $\nu$ $\zeta$
$0.5$ $3.17\pm 0.13$ $0.59\pm 0.04$ $0.63\pm 0.07$
$1.0$ $2.41\pm 0.06$ $0.59\pm 0.01$ $0.59\pm 0.04$
$2.0$ $0.41\pm 0.01$ $0.40\pm 0.01$ $0.49\pm 0.09$
$3.0$ $0.31\pm 0.03$ $0.62\pm 0.06$ $0.87\pm 0.05$
$4.0$ $0.17\pm 0.02$ $0.30\pm 0.05$ $0.32\pm 0.05$
Table 1: Critical disorder strength and exponents $\nu$, $\zeta$ for different chaos parameter, $k$. We have used $\langle J^{2}\rangle$ with $N=12,14$ and $16$ to identify the critical points with time averaging ranging from $n=10^{5}$ to $3\times 10^{5}$ for 100 different realizations.
Figure 6: (a) Plot of $\langle S_{N/2}(A)\rangle/N^{\zeta/\nu}$ vs $w$ for $k=1$ corresponding to different system sizes cross each other at $w=w_{c}\sim 1.8$. Inset shows the collapse of the data with $\nu=0.67$ and $\zeta=0.76$.
In summary, this work has shown the existence of a critical strength of permutation symmetry breaking, beyond which there is a second order phase transition from area law entangled states with total angular momentum scaling as $N$, to volume law entangled states with total angular momentum scaling as $\sqrt{N}$. The critical disorder strength shows clear dependence on the dynamics being regular or chaotic in the disorder-free permutation symmetric space. The exponents on the other hand seem to show a weaker dependence. Results not shown indicate that the criticality is robust on changing the the initial state except when it is in the regular region of the mixed phase space.
From an applications perspective, our results show that for disorder less than the critical strength, permutation symmetric systems are likely to be reliably used for their intended application, being at least not volume law entangled in the full Hilbert space. It is possible that the proposed phase transition be experimentally observed by extending few qubit experimental kicked top studies Chaudhury et al. (2009); Neill et al. (2016); Krithika et al. (2019), wherein disorder is anyway present. In fact, a disordered all-to-all interacting system, simulating the SYK model, has been recently experimentally studied using a combination of atomic cloud in a cavity and controllable light shift in the large detuning limit Sauerwein et al. (2023).
Among future directions, it is interesting to explore if breaking permutation symmetry through other methods, say by introducing a power law interaction such as in Russomanno et al. (2021), will also exhibit a similar phase transition as induced by disorder.
Acknowledgements: MC and UD acknowledege the HPC facility Chandra at IIT Palakkad where the computations were carried out. AL thanks Sylvia Pappalardi for discussions at an early stage of this work, at the program on “Dynamical Foundations of Many-Body Chaos”, Institut Pascal, Université Paris-Saclay, during March-April 2023.